# Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models

### v3

@@ 3.3 Matrix-Vector Method for Linear
In this section, we present a way to transform Linear operation into matrix-vector format. Figure 2 illustrates this process: Figure 2.a shows the linear transformation of multi-channel input: $\mathbf{W}\mathbf{x}=\mathbf{y}$. Figure 2.b provides a specific example of Figure 2.a, Figure 2.c converts the linear transformation in Figure 2.b into the corresponding matrix-vector representation: $\mathbf{W}^{\prime}\diamond\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$. Thus, the linear transformation can be represented in matrix-vector form as follows:
$\mathbf{x}_{i+1}=\mathbf{W}_{i}\mathbf{x}^{i}\rightarrow\mathbf{x}_{i+1}^{\prime}$ $=\mathbf{W}_{i}^{\prime}\diamond\mathbf{x}_{i}^{\prime}=(\mathbf{W}_{i}^{\prime})^{T}\mathbf{x}_{i}^{\prime}$ (5)
@@
$\rightarrow\mathbf{W}_{2}^{\prime}\sigma(\mathbf{W}_{1}^{\prime}\mathbf{x}^{\prime}+\mathbf{b}_{1}^{\prime})+\mathbf{b}_{2}^{\prime}$
Figure 3: The Transformer process.
Figure 4: The process of transforming $\hat{\mathbf{H}}=\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 5: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
@@ 3.4 Matrix-Vector Method for MHA
$\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},$ (8)
$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
@@
Our objective is to express the MHA as $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Prior to transforming the MHA into the matrix-vector form, we need to conduct a comprehensive analysis and clearly define the research object. In Eq. (7), $\hat{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation. The learning process for the input $\mathbf{x}$ primarily consists of two parts: $\mathbf{V}_{1},\ldots,\mathbf{V}_{h}$ and $\mathbf{H}_{1},\ldots,\mathbf{H}_{h}$. $\mathbf{H}_{i}$ is derived based on the parameters $\mathbf{x}_{i}$, $\mathbf{W}_{Qi}$ and $\mathbf{W}_{Ki}$.
Figure 4.a represents the computation process of $\mathbf{H}_{1}\mathbf{V}_{1}=\hat{\mathbf{H}}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}=\hat{\mathbf{H}}_{h}$. In Figure 4.b, we present a simple example of Figure 4.a. Figure 4.b depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}],i=1,2$. In Figure 4.c we convert Figure 4.b into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\mathbf{H}_{2}$ and $\mathbf{W}_{V1},\mathbf{W}_{V2}$. More details can be found in Appendix B.
Figure 5 illustrates the parameter transformation scenario after incorporating $\mathbf{W}_{O}$. Figure 5.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 5.b rewrites Figure 5.a as $\hat{\mathbf{H}}^{\prime}\diamond\mathbf{W}_{O}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Consequently, based on Figures 4 and 5, the entire MHA can be expressed as $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{HV}^{T}\diamond\mathbf{W}_{O}^{\prime}$ (Based on Figures 5, we can get $\mathbf{W}_{HVO}^{\prime}$, but the resulting expression is too complex and not easy to understand, so we did not compute the final result. However, it is easy to analyze that as long as the variables in $\mathbf{W}_{HV}^{T}$ and $\mathbf{W}_{O}^{\prime}$ are non-zero, the result is essentially a dense matrix). So the whole MHA can be written as:
@@
3.5 The UAT Format of Multi-Layer Transformer
The Transformer architecture is founded on two pivotal components: FFN and MHA. In Sections 3.3 and 3.4, we have showcased the matrix-vector representations for both FFN and MHA. In this section, we delve into why multi-layer Transformer is the implementation of UAT.
Based on Equations 6 and 10, we can demonstrate that $i+1$-th layer ($i=0,1,2,\ldots$) Transformer is a concrete realization of the UAT, which can be written as (see Appendix C for more details):
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (11)
where $\mathbf{x}_{i+1}$ is the output of ${i+1}$-th layer, $\mathbf{x}_{0}$ is the input of the network, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$. The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$ layer of UAT with $\mathbf{x}_{0}$ as input. This enhances the model’s ability to dynamically adjust functions based on input. In the MHA mechanism, the parameters change dynamically with the input. Therefore, in the formula above, all $\mathbf{W}^{\prime}_{j,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
@@ 4 Discussion
Leveraging on our proof in the previous section that the Transformer is the tangible embodiment of UAT, in this section we address the following critical problems and explore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming? What enables LLMs’ capacity for ICL? How does the LoRA scheme effectively fine-tune LLMs? What justifies the feasibility of pruning LLMs?
@@ 4.1 What Makes Transformer So Powerful in LLMs?
Theoretically, we have established that Transformer networks are concrete implementations of the UAT, enabling them to approximate any Borel measurable function. While UAT provides powerful function approximation capabilities, it inherently lacks the ability to approximate multiple functions simultaneously. However, language tasks are inherently diverse, often requiring the approximation of different functions based on the input. For instance, when summarizing, translating, or continuing the same text, the input functions are nearly identical, with only minor variations in the prompt. Without the ability to dynamically approximate functions based on the input, simply fitting a general function trend based on input will result in identical or similar outputs.
To address this, LLMs must distinguish and adapt to these nearly identical functions, dynamically generating response functions based on the input. The MHA and residual mechanisms in Transformers equip LLMs with the ability to dynamically approximate relevant functions according to the input. Specifically, MHA allows for the dynamic adjustment of the weight parameters in UAT in response to the input, while the residual mechanism supports the dynamic approximation of bias terms, as illustrated in Figure 6. This theoretical foundation enables Transformer-based LLMs to handle a wide range of tasks, including translation, continuation, summarization, code generation, and solving mathematical problems.
Figure 6: This diagram vividly demonstrates the output differences between an MLP and a Transformer when given the same input content (represented by the blue line in a) with only slight variations (represented by the black line in a). Since the input content is nearly identical and the MLP’s parameters are fixed, the MLP tends to produce the same output (b), influenced by the overall input function trend. However, the Transformer, with its ability to dynamically adapt to the input, generates different outputs based on the input variations. In the context of LLMs, the blue line in a can be understood as a consistent piece of text, while the black line represents different prompts, such as translation, continuation, or summarization.
@@ 4.2 What Enables LLMs to Possess ICL Capability?
So, how does this contextual interaction capability arise within LLMs? The formula $\mathring{\mathbf{H}}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ in Figure 5 reveals this mode of contextual interaction. Since $\mathbf{W}^{\prime}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts. This learning of holistic contextual information constitutes the foundation of contextual interaction within LLMs. (See Appendix D for more details)
@@ 4.3 What Justifies the Feasibility of Pruning LLMs?
$|\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|<\varepsilon\quad.$ (12)
for all $\mathbf{x}\in\mathbf{I}_{n}$. Let’s assume $\Lambda=\{1,2...N\},\Lambda_{1}\cup\Lambda_{2}=\Lambda,\Lambda_{1}\cap\Lambda_{2}=\emptyset$ and $|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$. Then we have:
$|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (13)
$+$ $\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|<\varepsilon$
Since $|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$, we have the following inequality:
@@
$|\sum_{j=1}^{N_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|$ (15)
$<\varepsilon+|\sum_{j=1}^{N_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|$

### v4

@@ 3.1 Matrix-Vector Method for Linear
In this section, we present a way to transform Linear operation into matrix-vector format. Figure 3 illustrates this process: Figure 3.a shows the linear transformation of multi-channel input: $\mathbf{W}\mathbf{x}=\mathbf{y}$. Figure 3.b provides a specific example of Figure 3.a, Figure 3.c converts the linear transformation in Figure 3.b into the corresponding matrix-vector representation: $\mathbf{W}^{\prime}\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$. Thus, the linear transformation can be represented in matrix-vector form as follows:
$\mathbf{x}_{i+1}=\mathbf{W}_{i}\mathbf{x}_{i}\mapsto\mathbf{x}_{i+1}^{\prime}=\mathbf{W}^{\prime}_{i}\mathbf{x}^{\prime}_{i}$ (5)
@@
$\mapsto\mathbf{W}_{2}^{\prime}\sigma(\mathbf{W}_{1}^{\prime}\mathbf{x}^{\prime}+\mathbf{b}_{1}^{\prime})+\mathbf{b}_{2}^{\prime}$
@@ 3.2 Matrix-Vector Method for MHA
$\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})$ (8)
$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
@@
Figure 4: The Transformer process.
Our objective is to express the MHA as $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Prior to transforming the MHA into the matrix-vector form, we need to conduct a comprehensive analysis and clearly define the research object. In Eq. 7, $\hat{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation. The learning process for the input $\mathbf{x}$ primarily consists of two parts: $\mathbf{V}_{1},\ldots,\mathbf{V}_{h}$ and $\mathbf{H}_{1},\ldots,\mathbf{H}_{h}$. $\mathbf{H}_{i}$ is derived based on the parameters $\mathbf{x}_{i}$, $\mathbf{W}_{Qi}$ and $\mathbf{W}_{Ki}$.
Figure 5.a represents the computation process of $\mathbf{H}_{1}\mathbf{V}_{1}=\hat{\mathbf{H}}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}=\hat{\mathbf{H}}_{h}$. In Figure 5.b, we present a simple example of Figure 5.a. Figure 5.b depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}],i=1,2$. In Figure 5.c we convert Figure 5.b into the matrix-vector form $(\mathbf{W}_{HV}^{\prime})^{T}\mathbf{x}^{\prime}=(\hat{\mathbf{H}}^{\prime})^{T}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\mathbf{H}_{2}$ and $\mathbf{W}_{V1},\mathbf{W}_{V2}$. More details can be found in Appendix B.
Figure 5: The process of transforming $\hat{\mathbf{H}}=\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $(\mathbf{W}_{HV}^{\prime})^{T}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 6 illustrates the parameter transformation scenario after incorporating $\mathbf{W}_{O}$. Figure 6.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 6.b rewrites Figure 6.a as $(\mathbf{W}_{O}^{\prime})^{T}\hat{\mathbf{H}}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Consequently, based on Figures 5 and 6, the entire MHA can be expressed as $(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{O}^{\prime}\mathbf{W}_{HV}^{\prime}$ (Based on Figures 6, we can get $\mathbf{W}_{HVO}^{\prime}$, but the resulting expression is too complex and not easy to understand, so we did not compute the final result. However, it is easy to analyze that as long as the variables in $\mathbf{W}_{HV}$ and $\mathbf{W}_{O}^{\prime}$ are non-zero, the result is essentially a dense matrix). So the whole MHA can be written as:
@@
Figure 6: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
3.3 The DUAT Format of Multi-Layer Transformer
The Transformer architecture is founded on two pivotal components: FFN and MHA. In Sections 3.1 and 3.2, we have showcased the matrix-vector representations for both FFN and MHA. In this section, we delve into why multi-layer Transformer is the implementation of DUAT.
Based on Eq 6 and 28, we can demonstrate that $i+1$-th layer ($i=0,1,2,\ldots$) Transformer can be written as (see Appendix C for more details):
$\mathbf{x}_{i+1}=$ $(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})$ (11)
$+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
where $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$. We define the parameters for cases where $i\geq 1$. For each $j=1,2,\ldots,i$, the updates are as follows: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,3}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}^{\prime}_{j,3}$. Additionally, for $j=2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (12)
$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$ layer UAT with $\mathbf{x}_{0}$ as input. In the MHA mechanism, the parameters change dynamically with the input. Therefore, in the formula above, all $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, are dynamically adjusted based on the input. According to Eq. 11, we know the mathematical format of a multi-layer Transformer is the same as UAT, but the parameters in the multi-layer Transformer’s UAT are not fixed, they will change with input. So we call the UAT format of parameters dynamically adjust based on input as DUAT.
@@ 4 Discussion
Leveraging on our proof in the previous section that the Transformer is the tangible embodiment of DUAT, in this section we address the following critical problems and explore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming? What enables LLMs’ capacity for ICL? How does the LoRA scheme effectively fine-tune LLMs? What justifies the feasibility of pruning LLMs?
@@ 4.1 What Makes Transformer So Powerful in LLMs?
Theoretically, we have established that Transformer networks are concrete implementations of the DUAT, enabling them more powerful approximation ability than UAT. While UAT provides powerful function approximation capabilities, it inherently lacks the ability to approximate multiple functions simultaneously. However, language tasks are inherently diverse, often requiring the approximation of different functions based on the input. For instance, when summarizing, translating, or continuing the same text, the input functions are nearly identical, with only minor variations in the prompt. Without the ability to dynamically approximate functions based on the input, simply fitting a general function trend based on input will result in identical or similar outputs.
To address this, LLMs must distinguish and adapt to these nearly identical functions, dynamically generating response functions based on the input. The MHA and residual mechanisms in Transformers equip LLMs with the ability to dynamically approximate relevant functions according to the input. Specifically, MHA allows for the dynamic adjustment of the weight parameters in UAT in response to the input, while the residual mechanism supports the dynamic approximation of bias terms. This theoretical foundation enables Transformer-based LLMs to handle a wide range of tasks, including translation, continuation, summarization, code generation, and solving mathematical problems.
@@ 4.2 What Enables LLMs to Possess ICL Capability?
So, how does this contextual interaction capability arise within LLMs? The formula $\mathring{\mathbf{H}}=(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}$ in Figure 6 reveals this mode of contextual interaction. Since $\mathbf{W}^{\prime}_{HVO}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts. This learning of holistic contextual information constitutes the foundation of contextual interaction within LLMs. (See Appendix D for more details)
@@ 4.3 What Justifies the Feasibility of Pruning LLMs?
$|\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+b_{j}\right)-f(\mathbf{x})|<\varepsilon\quad.$ (13)
for all $\mathbf{x}\in\mathbf{I}_{n}$. Let’s assume $\Lambda=\{1,2...N\},\Lambda_{1}\cup\Lambda_{2}=\Lambda,\Lambda_{1}\cap\Lambda_{2}=\emptyset$ and $|\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$. Then we have:
$|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+b_{j}\right)$ (14)
$+$ $\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+b_{j}\right)-f(\mathbf{x})|<\varepsilon$
Since $|\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$, we have the following inequality:
@@
$|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+b_{j}\right)-f(\mathbf{x})\mid$ (16)
$<$ $\varepsilon+|\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+b_{j}\right)|$